# X-ray2CTPA: Generating 3D CTPA scans from 2D X-ray conditioning

### v1

@@ Introduction
CXR to CT generation. CXR to CT conversion is an under-explored area of research and there are not many works that attempt to solve this problem. X2CT-GAN [8] explores the use of Generative Adversarial Networks (GANs) for reconstructing 3D CT volumes from biplanar 2D CXRs. MedNeRF [9] introduces a deep learning model for reconstructing CT projections from as few as a single-view CXR. This approach utilizes a neural radiance field (NeRF) architecture to learn a continuous representation of CT scans. The model effectively disentangles the shape and volumetric depth of surface and internal anatomical structures from 2D images. For both papers, the CXRs used are synthetic constructions from CTs called DRR, which are similar but not identical to real CXRs. This is due to the lack of a publicly available dataset of paired CXR and CT. We argue that the task of converting DRRs to CT scans is less complex than our task because DRRs inherently possess more information than CXRs. Also, DRRs are naturally aligned and registered with CTPAs due to the way they are constructed.
@@ Results
Figure 2: Qualitative results. We show examples of CXR to 3D CTPA Generation with an X-ray2CTPA model. (A) Shown is a generated 3D-CTPA sample from the test set, generated from a 2D CXR input. We compare slices from the generated 3D-CTPA scan to the ground truth sample. For simplicity, we present every eighth slice from the whole 3D-scan. (B) The input CXR along with its matching coronal slices from the matching CT and generated 3D-CTPA. This is the same sample presented in (A). Note: Throughout the paper the axial CT images appear slightly deformed. This is due to the preprocessing of the data and region of interest selection. (We elaborate on this in the Dataset Section.)
@@
In Fig. 2 we show visual examples of generated scans. The results demonstrate that the generated scans seem anatomically correct, realistic and consistent between sequential slices. Fig. 2. (A) visualizes slices from a generated sample from our held-out test set and the corresponding slices compared with the ground truth scan. As can be seen, the slices are similar but variations exist compared to the ground truth sample. These differences will be further discussed in the following sections. In Fig. 2. (B) the input CXR for the generated sample is depicted, along with the corresponding coronal example slice from the ground truth and the matching slice from the generated sample. It is evident that general anatomical structure is preserved between the generated sample and ground truth and matches the CXR input.
@@
Table 2 evaluates several design choices and the contribution of different network components. We examined different loss functions as alternatives for the model training. In addition to $L_{1}$ loss which minimizes the pixel distance between the generated sample and the ground truth, we compared the use of adding perceptual loss [23], adding adversarial loss using a discriminator and classification loss. As can be seen from the results, adding the perceptual loss improved the results compared to using only $L_{1}$ loss. However, while the addition of adversarial and classification loss did not enhance the quantitative metrics, they both played important roles for our model. Adding the adversarial loss resulted in more visually realistic samples. The classification loss improved the classification results. This is presented in Fig. 4.
@@
L1 + LPIPS $\textbf{1.89}\pm\textbf{0.64}$ $\textbf{14.78}\pm\textbf{0.60}$ $\textbf{0.22}\pm\textbf{0.07}$ $\textbf{0.34}\pm\textbf{0.06}$
@@
X-ray2CTPA $1.94\pm 0.64$ $13.89\pm 0.75$ $0.22\pm 0.06$ $0.35\pm 0.06$
In addition to the quantitative results, we also compare the visual differences of using different loss functions (Fig. 4). All loss functions were trained as a weighted combination with the $l1$ loss. In general, it is evident that the generated samples are like the ground truth, although not identical. Supporting the quantitative results, the LPIPS loss improves both the visual appearances of the generated samples and seem to guide the scans to be closer in structure and color contrast to the ground truth. The classification loss adds distortion to the samples, which is expected as its purpose is only to aid the classification task. The adversarial loss provides very realistic images but, in some cases, the generated sample is derived further away from the ground truth. The weighted combination of all losses (Fig. 4 - "Full") provides the best tradeoff between all losses.
@@
As expected, the real CTPA classification results outperform the CXR results achieving an AUC of 0.827 [95% CI: 0.75-0.92] compared to 0.691 [95% CI: 0.54-0.77]. The classification using the generated CTPAs significantly outperformed the CXR classification results with an AUC of 0.766 [95% CI: 0.7-0.89]. In addition, the specificity and NPV of the model improve as well compared to the CXR classifier. In a clinical context, this outcome suggests a higher number of true negative results, leading to a reduction in unnecessary CTPA scans performed. It is evident that using the synthetic CTPAs enhance performance compared to using solely CXRs. For the CTPAs classification, we used the latent output representation from the 2D-VAE used in our diffusion model. The CXRs were classified using RadImageNet [26], a ResNet-50 based model pretrained on millions of radiologic images.
@@
Generated CTPA only classifier $0.766$ [$0.7$-$0.89$] ±$\pm 0.07$ $73.17$ $72.88$ $73.91$ $51.52$ $87.76$
@@
The quantitative generation results for the LIDC dataset are presented in Table 4. As can be seen, although trained on an extremely low number of samples - the evaluation metrics outperform the CTPA generation results.
@@
While X-ray2CTPA model can generate high-quality 3D CTPA samples that closely resemble the ground truth in structure and appearance, there are instances where it diverges significantly or fails to accurately replicate the ground truth. Some examples from our test set as presented in Fig. 6. The main cases where the model fails are when rare abnormalities or additional pathologies (other than PE) are presented. In some of these cases the model generates the pathology in some variant but may replicate it in a wrong location or fail to generate the pathology whatsoever. We hypothesis that this issue may be caused by the high variability nature of the CTPA samples and due to the fact, that in many cases PE is secondary to other patient co-morbidities. We believe that training the model with a larger dataset may solve this issue in future work.
Figure 6: Failed Examples and Hallucinations. Presented are example slices from our test set where we consider the model to fail. (A) Moderate pleural effusions and kyphoscoliosis pathologiesare apparent in both the CXR input and the ground truth sample. The generated sample shows small pleural effusions and "straightened" back positioning. (B) A lung cancer in the left lung is absent in the the generated CTPA. (C) Hallucinates the presence of pleural effusions. (D) Hallucinates the presence of a consolidation in the left lung. (E) Deformed right chest wall is absent.
@@ Discussion
We were able to generate somewhat realistic, consistent and anatomically correct 3D scans. The generated 3D scans are not identical to the ground truth CTPAs but are similar and capture their structure. We experimented in the image size and additional modifications to the model and loss functions. Our best performing model improved the CXR PE classification results by 7%; Improving from an AUC of 0.691 [95% CI: 0.54-0.77] to 0.766 [95% CI: 0.7-0.89]. In addition, the specificity of the model improved, which in a clinical setting may potentially lead to a reduction of redundant CTPA scans performed for PE suspected patients. Unlike other diffusion studies in the medical field, this research uniquely applies generated data to address a real-life clinical challenge rather than focusing on the quality of the generated samples. As far as we know, this is the first model that attempts to convert 2D chest X-rays to CTPA scans. We also demonstrate the model’s generalizability by fine-tuning it on an extremely small number of samples to achieve 2D CXR to 3D CT translation.
@@ Methods
All models were trained on a single NVIDIA RTX A5000 GPU, and each took about a week of training. The code for our model was based on Medical Diffusion [11] with various alterations. The training parameters included a base learning rate of $1e^{-4}$ for the unconditional setting and when pretraining using RSPECT dataset and a learning rate of $1e^{-5}$ when fine-tuned for the final conditional model. All were trained with a batch size of 2 with Adam optimizer and Cosine Anneal scheduler. We used DDPM sampling with 1000 steps.

### v3

@@ Introduction
CXR to CT generation. CXR to CT conversion is an under-explored area of research and there are not many works that attempt to solve this problem. X2CT-GAN [8] and it’s extension - XprospeCT[16], explores the use of Generative Adversarial Networks (GANs) for reconstructing 3D CT volumes from biplanar 2D CXRs. MedNeRF [9] introduces a deep learning model for reconstructing CT projections from as few as a single-view CXR. This approach utilizes a neural radiance field (NeRF) architecture to learn a continuous representation of CT scans. The model effectively disentangles the shape and volumetric depth of surface and internal anatomical structures from 2D images. For both papers, the CXRs used are synthetic constructions from CTs called DRR, which are similar but not identical to real CXRs. This is due to the lack of a publicly available dataset of paired CXR and CT. We argue that the task of converting DRRs to CT scans is less complex than our task because DRRs inherently possess more information than CXRs. Also, DRRs are naturally aligned and registered with CTPAs due to the way they are constructed.
@@ Results
Figure 2: Qualitative results. We show examples of CXR to 3D CTPA Generation with an X-ray2CTPA model. Shown are four randomly selected generated 3D-CTPA samples from the test set, synthesized from a 2D CXR input. We compare slices from the generated 3D-CTPA scan to the ground truth samples. For simplicity, we present five key slices from the whole 3D-scan. The input CXR is presented as well, along with its matching coronal slice from the matching CTPA and generated 3D-CTPA. Note: throughout the paper the axial CT images appear slightly deformed. This is due to the preprocessing of the data and region of interest selection. (We elaborate on this in the Dataset Section.)
@@
In Fig. 2 we show visual examples of generated scans. The results demonstrate that the generated scans seem anatomically correct, realistic and consistent between sequential slices. Fig. 2. visualizes key slices from four randomly selected generated samples from our held-out test set and the corresponding slices compared with the ground truth scan. As can be seen, the slices are similar but small variations exist compared to the ground truth samples. These differences will be further discussed in the following sections. In addition, the input CXRs for the generated samples are depicted, along with the corresponding coronal example slice from the ground truth and the matching slice from the generated sample. It is evident that general anatomical structure is preserved between the generated sample and ground truth and matches the CXR input.
@@
Table 2 evaluates several design choices and the contribution of different network components. We examined different loss functions as alternatives for the model training. In addition to $L_{1}$ loss which minimizes the pixel distance between the generated sample and the ground truth, we compared the use of adding perceptual loss [24], adding adversarial loss using a discriminator and classification loss. As can be seen from the results, adding the perceptual loss improved the results compared to using only $L_{1}$ loss. However, while the addition of adversarial and classification loss did not enhance the quantitative metrics, they both played important roles for our model. Adding the adversarial loss resulted in more visually realistic samples. The classification loss improved the classification results. This is presented in Fig. 4. Additionally, we modified the pretraining process to include the RSPECT dataset not only in as an unconditional dataset for CTPA scans but also in a conditional setting by incorporating synthetic CXR images or digitally reconstructed Radiographs (DRRs) created from the CTPA scans. Adding this pretraining step highly boosted our performance by all metrics.
@@
L1 + LPIPS $1.89\pm 0.64$ $14.78\pm 0.60$ $0.22\pm 0.07$ $0.34\pm 0.06$
@@
X-ray2CTPA $\textbf{0.29}\pm\textbf{0.22}$ $\textbf{19.41}\pm\textbf{0.96}$ $\textbf{0.58}\pm\textbf{0.08}$ $\textbf{0.14}\pm\textbf{0.04}$
In addition to the quantitative results, we also compare the visual differences of using different loss functions (Fig. 4). All loss functions were trained as a weighted combination with the $l1$ loss. In general, it is evident that the generated samples are like the ground truth, although not identical. Supporting the quantitative results, the LPIPS loss improves both the visual appearances of the generated samples and seem to guide the scans to be closer in structure and color contrast to the ground truth. The pretraining step using the DRRs as conditioning further enhances the LPIPS loss improvements and provides the most realistic images which are also the most close to the ground truth. The classification loss adds distortion to the samples, which is expected as its purpose is only to aid the classification task. The adversarial loss provides very realistic images but, in some cases, the generated sample is derived further away from the ground truth. The weighted combination of all losses (Fig. 4 - "Full") provides the best tradeoff between all losses.
@@
As expected, the real CTPA classification results outperform the CXR results achieving an AUC of 0.827 [95% CI: 0.75-0.92] compared to 0.691 [95% CI: 0.54-0.77]. The classification using the generated CTPAs significantly outperformed the CXR classification results with an AUC of 0.803 [95% CI: 0.72-0.9]. In addition, the specificity and NPV of the model improve as well compared to the CXR classifier. In a clinical context, this outcome suggests a higher number of true negative results, leading to a reduction in unnecessary CTPA scans performed. It is evident that using the synthetic CTPAs enhance performance compared to using solely CXRs. For the CTPAs classification, we used the latent output representation from the 2D-VAE used in our diffusion model. The CXRs were classified using RadImageNet [27], a ResNet-50 based model pretrained on millions of radiologic images.
@@
Generated CTPA only classifier $0.80$ [$0.72$-$0.9$] ±$\pm 0.07$ $74.0$ $74.1$ $73.91$ $54.8$ $87.0$
@@
The quantitative generation results for the LIDC dataset are presented in Table 4. As can be seen, although trained on an extremely low number of samples - the evaluation results are high.
@@
While X-ray2CTPA model can generate high-quality 3D CTPA samples that closely resemble the ground truth in structure and appearance, there are instances where it slightly diverges or fails to accurately replicate the ground truth. Some examples from our test set as presented in Fig. 6. The main areas where the model fails are around small veins or when rare abnormalities or additional pathologies (other than PE) are presented. In some of these cases the model generates the pathology in some variant but may replicate it in a wrong location or fail to generate the pathology whatsoever. We hypothesis that this issue may be caused by the high variability nature of the CTPA samples and due to the fact, that in many cases PE is secondary to other patient co-morbidities. We believe that training the model with a larger dataset may solve this issue in future work.
Figure 6: Failed Examples and Hallucinations. Presented are example slices from our test set where we consider the model to fail. (A) Moderate pleural effusions and kyphoscoliosis pathologiesare apparent in both the CXR input and the ground truth sample. The generated sample shows the pleural effusions but they are not generated to be identical to the ground truth. (B) A lung cancer in the left lung is generated in the synthesized CTPA at the correct location and but is slightly different in texture and color. (C,D) Hallucinates the shape and locations of veins around the heart area. (E) Deformed right chest wall is absent.
@@ Discussion
We were able to generate somewhat realistic, consistent and anatomically correct 3D scans. The generated 3D scans are not identical to the ground truth CTPAs but are similar and capture their structure. We experimented in the image size and additional modifications to the model and loss functions. Our best performing model improved the CXR PE classification results by 11%; Improving from an AUC of 0.691 [95% CI: 0.54-0.77] to 0.803 [95% CI: 0.72-0.9]. In addition, the specificity of the model improved, which in a clinical setting may potentially lead to a reduction of redundant CTPA scans performed for PE suspected patients. Unlike other diffusion studies in the medical field, this research uniquely applies generated data to address a real-life clinical challenge rather than focusing on the quality of the generated samples. As far as we know, this is the first model that attempts to convert 2D chest X-rays to CTPA scans. We also demonstrate the model’s generalizability by fine-tuning it on an extremely small number of samples to achieve 2D CXR to 3D CT translation.
@@ Methods
All models were trained on a single NVIDIA RTX A5000 GPU, and each took about a week of training. The code for our model was based on Medical Diffusion [11] with various alterations. The training parameters included a base learning rate of $1e^{-4}$ for the unconditional setting and when pretraining using RSPECT dataset and a learning rate of $1e^{-5}$ when fine-tuned for the final conditional model. An additional pretraining step was added in which we used the RSPECT dataset for pretraining but in a conditional setting. We used artificial CXR images or digitally reconstructed Radiographs (DRRs) as conditioning. The DRRs were created using DiffDRR [35] Python package for differentiable X-ray rendering. All were trained with a batch size of 2 with AdamW optimizer and Cosine Anneal scheduler. We used DDPM sampling with 1000 steps.